REAL-TIME L^{A}TEX DEMONSTRATION MARCH 2, 1998

DOUG MEADE AND GEORGE MCNULTY

Here is some emphasized text. And, now, some mathematics α that is in a paragraph. β Hereissometexttypesetinmathmode. And, again, with displayed math mode:

αβAnd, asbefore, sometext.

But, this time it is not a part of the paragraph.

And, as before, some text. the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor.

the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor. the, if, floor.

Here are some math symbols that I know:

⟶⟹⟵⟸ ⋄♡
Font  Example
regular  x
mathit  x
mathrm  x
mathbf  x
mathsf  x
mathtt  x
regular  X
mathbb  𝕏
mathcal  X
regular  +
boldsymbol  +
mathbf  +
(\begin{matrix}0 & 1 \\ 2 & 3 \end{matrix})

\begin{cases}0 & 1 & 2 \\ 3 & 4 & x^{2} \end{cases}∥

Here is a matrix that is typeset in a paragraph. | \begin{matrix}0 & 1 \\ 2 & 3 \end{matrix} | = det ( \begin{matrix}0 & 1 \\ 2 & 3 \end{matrix} ) See, isn’t that neat?

Date: March 2, 1998.

Next, we look at fractions: \frac{α}{1+∞} and roots \sqrt{β} and \sqrt[3]{1 + \frac{1+x}{1+α}} Superscipts and subscripts: σ_{3} and σ^{3}. What will \sigma^3.14 yield? Answer: σ^{3}.14.

α_{γ_{ϵ}^{δ}}^{β}

Sums, etc.: inline: ∑_{i=0}^{i=100} displayed:

∑_{i=0}^{i=100}